# LaMSUM: Creating Extractive Summaries of User Generated Content using LLMs

### v2

@@ 5.4. What Fails to Deliver Results?
To ensure extractive summarization, we tested an additional approach – each sentence is tagged with a sentence number, LLM is prompted to select the best $q$ sentences and output only the sentence numbers of the best $q$ sentences. Thereafter, the sentences corresponding to the sentence numbers can be retrieved. For instance, if $s$ is 200 and $q$ is 100, the task is to output the sentence numbers of the best 100 sentences from a pool of 200 sentences. In such cases, LLMs hallucinate and provide an output consisting of either all the odd number sentences or all the even number sentences.
@@
6. Ensembling of LLMs in LaMSUM
Figure 4. Ensembling applied at chunk using all the four models. $q$ length summaries are obtained from each chunk through different LLMs using the shuffling method discussed in Section 3.3. All the summaries are ensembled through voting algorithms to achieve the final summary of length $q$. This final $q$ length summary is thereafter forwarded to the next level.
 Ensemble Voting Method R1 R2 RLSum
 All 4 models Plurality 54.16 20.79 53.00
Proportional 54.70 19.95 53.14
Ranked 51.78 16.97 50.24
Best 3 models Plurality 58.88 21.74 57.60
Proportional 58.04 22.78 56.71
Ranked 57.22 19.51 55.33
Weak 3 models Plurality 50.64 18.82 49.31
Proportional 52.92 19.23 51.07
Ranked 47.50 16.74 46.37
Table 4. Results obtained through ensembling of LLMs in LaMSUM for Claritin dataset. We compare 3 different cases: i) ensemble of all 4 models – Llama, Mixtral, Gemini and GPT-4o mini ii). ensemble of best 3 models – Llama, Mixtral and GPT-4o mini iii). ensemble of weak 3 models – Llama, Mixtral and Gemini.
Different LLMs can be considered as distinct experts, each offering their vote or opinion about the textual unit to be included in the summary. In LaMSUM, each expert (LLM) operated independently, determining which sentences were worthy of inclusion. We hypothesize that using an ensemble of LLMs in LaMSUM, where these experts collaborate and reach a consensus, has potential to improve performance of the overall system (Verga et al., 2024; Schoenegger et al., 2024).
We apply ensembling at each chunk as shown in Figure 4. For a given chunk $i$ containing $s$ textual units, the goal is to produce a final summary of size $q$ by gathering opinions of all the experts. Ensembling takes place in two stages. Stage 1 - For each chunk we obtain $q$ length summary through different LLMs by using shuffling and voting method as demonstrated in Section 3.3. Stage 2 - To ensemble the $q$ length summaries from different LLMs, we again leverage the voting methods. In stage 2 we make use of the same voting method that was used in stage 1 to maintain consistency. The summary generated from the ensembling process is treated as the final summary for chunk $i$ and is carried forward to the next level.
As shown in Table 3, for Claritin dataset, GPT-4o mini outperforms the other models, while Gemini delivers weaker results. We evaluate three different scenarios based on the number of LLMs used in the ensembling process: i) all four models – Llama, Mixtral, Gemini and GPT-4o mini, ii) the top three models – Llama, Mixtral and GPT-4o mini, iii) the three weaker models – Llama, Mixtral and Gemini. Table 4 shows the results of ensembling in LaMSUM for different scenarios and voting methods.
In the approval-based voting algorithm, using all four models or just the top three models for ensembling doesn’t significantly improve performance; the resulting performance after ensembling remains close to that of the weakest model in each case. However, ensembling with the ranked-based voting algorithm yielded interesting results. When ranked voting is applied to individual LLMs, it delivers the worst performance compared to other voting algorithms, as shown in Table 3. When using ranked voting with all four models in the ensembled LaMSUM, the performance is close to that of the weakest model – Gemini (49.39) ¡ Llama (49.61) ¡ Ensemble (50.24) ¡ Mixtral (54.65) ¡ GPT-4o mini (56.05). However, when considering the ensembled LaMSUM with the top three models – Gemini (49.39) ¡ Llama (49.61) ¡ Mixtral (54.65) ¡ Ensemble (55.33) ¡ GPT-4o mini(56.05), ensemble outperforms Mixtral, and the ROUGE score increases from 50.24 to 55.33. While the ensembled LaMSUM with the top three models does not surpass GPT-4o mini, it does enhance the performance of the overall setup.
Takeaway: Ensemble of LLMs in LaMSUM yield good results with rank-based voting method, and can help improve the performance of the overall setup if we select models for ensembling that perform well individually in LaMSUM. Including all models in the ensemble could result in the weakest model becoming the bottleneck.
7. Does Fine-Tuning Help?
The performance of LLMs on specific downstream tasks, such as extractive summarization, can be suboptimal when not explicitly fine-tuned. LLMs tend to generate erroneous information while sounding persuasive and assured (Bang et al., 2023; Xu et al., 2024). In Section 2, we highlighted instances where LLMs exhibit nonplussed responses by generating abstractive summaries despite being prompted for extractive summarization. They also alter some words, so we perform output calibration in LaMSUM. To address these challenges, we explore the efficacy of instruction fine-tuning through explicit instructions (Wei et al., 2022) and compare it with LaMSUM.
We utilize the Claritin dataset for instruction fine-tuning as it has the highest number of tweets among all the three datasets. It is divided into training (60%), validation (10%), and test (30%) set. The US-Election and MeToo datasets are retained as additional test sets for evaluation. Instruction fine-tuning involves three key components: i) Instruction: The prompts used for zero-shot prompting, as detailed in Section 3.3.2. ii) Input: Set of input text $\mathcal{T}$ to be summarized. iii) Output: The reference summary comprises sentences extracted from $\mathcal{T}$.
We fine-tune an open-source LLM, llama3.1-8b from Meta (Dubey and et al., 2024), using a 15 GB Tesla T4 GPU. During the fine-tuning phase, we set the learning rate, warmup ratio, and batch size to 2e-4, 0.0, and 2, respectively. For this process, we employ PEFT (Parameter Efficient Fine Tuning), which involves freezing the layers of the pre-trained model and only fine-tuning the last few layers specific to the downstream task. We follow the QLoRA (Dettmers et al., 2023) method in PEFT training. QLoRA is an optimized variant of LoRA (Hu et al., 2021) that reduces the precision of weight parameters to 4-bit. This reduction in precision decreases the model size, which is advantageous in scenarios with limited memory available for fine-tuning.
Table 5 shows the results of the fine-tuned llama3.1-8b model. We observed that the outputs are robust and adhere closely to the instructions provided. The generated summaries contain the exact sentences extracted from the input text, but the performance of the fine-tuned model is inferior compared to LaMSUM.
 Dataset R1 R2 RLSum
 Claritin 18.60 3.11 17.14
US-Election 26.55 14.81 25.05
MeToo 28.94 10.81 24.80
Table 5. Metric scores from fine-tuned llama3.1-8b for various datasets.
Takeaway: Fine-tuning an LLM with the appropriate instructions, input, and output can facilitate extractive summarization. But fine-tuning requires a substantial amount of data with gold-standard summaries. With less data, the model may overfit, resulting in poor performance, as shown in the Table 5. Additionally, a model fine-tuned on one dataset may not perform well on another dataset. Our proposed LaMSUM addresses these issues by being effective when the input data is large enough to exceed a single context window but small enough to be used for fine-tuning.
8. Concluding Discussion
This work marks an early attempt to achieve extractive summarization of large user-generated text that exceeds a single context window using zero-shot learning. The proposed multi-level framework LaMSUM leverages approval based and ranked based voting algorithms to generate robust summaries. Experiments conducted on three distinct crowd-sourced datasets demonstrated the efficacy of LaMSUM, as it outperformed the results achieved by state-of-the-art fine-tuned models. Ensembling of LLMs in our proposed framework can enhance the performance of weaker models.
Note that there can be a concern regarding the potential data leakage, as the experiments involve newer LLMs that may have been exposed to the experimented datasets during their pre-training phase. In Section 5.2 we showcased that the vanilla LLM, which also includes LLM underperformed, whereas our proposed framework which generates robust summaries yielded good results. This highlights the efficacy of our model, even when it is exposed to data leakage.
Limitation
Our proposed framework, LaMSUM, very well handles text of any length, conditioned on the fact that the final summary fits within a single context window. Some modifications to LaMSUM may be necessary when the output summary exceeds the size of a single context window. We focused only on the user-generated text where each textual unit is independent of the others; future research could extend the framework to summarization tasks involving contextual dependency such as book summarization.
Ethical Considerations
Our research focuses on using LLMs to produce extractive summaries for user-generated text. Given the recent rise of LLMs and the growing interest in applying them across different research fields, we believe this research direction can help unveil the potential of LLMs. Below, we outline concerns that need to be considered and addressed in this research area:
  • Bias and discrimination: LLMs are biased towards their training data, which can lead them to favor certain textual units when creating summaries.
  • Transparency and accountability: LLMs are black boxes, with an opaque decision-making process, making it difficult to discern how or why specific textual units are chosen for summarization. This lack of transparency can create ethical issues, especially in scenarios where clear explanations for decisions are needed.
  • Environmental impact: The training and operation of LLMs require substantial computational resources, which contribute to increased energy consumption and may have a negative environmental impact.
While LLMs can produce high-quality summaries, their use must be approached with careful consideration of potential ethical implications.

### v4

@@ Appendix A Appendix
To ensure extractive summarization, we tested an additional approach – each sentence is tagged with a sentence number, LLM is prompted to select the best $q$ sentences and output only the sentence numbers of the best $q$ sentences. Thereafter, the sentences corresponding to the sentence numbers can be retrieved. For instance, if $s$ is 100 and $q$ is 50, the task is to output the sentence numbers of the best 50 sentences from a pool of 100 sentences. In such cases, LLMs hallucinate and provide an output consisting of either all the odd number sentences or all the even number sentences.
@@
Safe City Post Features
  Feature Details
 id unique id for each post
lang_id language id
building building where incident took place
landmark landmark near the place of incident
area area where incident occurred
city city where incident happened
state name of the state
country country name
latitude coordinates information
longitude coordinates information
created_on date when the post is made
description details about the incident
additional_detail more information about the incident
age age of the person
gender_id gender id
gender gender of the person
incident_date date when the incident took place
is_date_estimate binary value – yes or no
time_from start time of the incident
time_to end time of the incident
is_time_estimate binary value – yes or no
categories harassment category
Table 8: Features or attributes associated with a post in Safe City platform. For our task, we utilise only the description feature.
  City A City B City C City D City E
Models R1 R2 RLSum R1 R2 RLSum R1 R2 RLSum R1 R2 RLSum R1 R2 RLSum
                 Vanilla LLM
 Llama 56.506 25.493 54.098 48.798 25.393 47.674 53.321 24.494 51.082 49.976 25.516 47.848 51.713 23.452 49.015
Mixtral 53.009 21.994 49.848 42.948 19.178 41.970 52.337 24.547 49.392 48.066 26.513 46.000 52.250 24.946 50.195
GPT 57.666 25.543 55.044 46.518 24.491 44.952 56.255 26.828 54.140 52.144 29.078 50.136 50.199 23.773 47.840
                 LaMSUM + Plurality Voting
 Llama 57.679 26.446 55.275 45.990 19.033 44.249 54.789 25.282 52.364 50.119 23.849 47.773 52.118 22.047 49.264
Mixtral 56.757 24.740 53.805 43.639 18.116 41.414 53.511 23.155 50.916 52.277 26.102 48.121 53.538 25.938 50.562
GPT 58.140 28.705 55.733 53.471 26.895 52.017 54.684 22.729 51.916 52.340 26.938 50.102 54.315 27.655 51.943
                 LaMSUM + Proportional Voting
 Llama 58.392 27.108 55.665 50.649 23.808 48.517 57.137 25.105 54.678 52.129 26.270 49.619 53.800 26.845 50.691
Mixtral 57.175 27.634 54.600 47.158 23.707 45.989 54.314 26.109 51.072 54.463 30.333 52.058 55.656 27.438 53.134
GPT 60.440 30.006 58.355 56.296 32.109 55.229 58.708 29.022 56.428 54.855 28.124 52.504 55.921 29.919 53.426
                 LaMSUM + Borda Count
 Llama 58.098 26.251 56.092 49.007 23.377 47.612 55.296 24.912 52.637 50.295 26.034 48.265 52.582 25.136 49.978
Mixtral 56.681 25.156 53.525 46.293 21.928 44.820 52.980 19.953 49.414 53.722 27.817 50.869 53.838 25.767 51.680
GPT 59.188 27.080 56.585 49.107 24.114 47.250 54.167 22.718 51.847 51.858 27.812 49.761 51.226 24.749 48.956
Table 9: Table showing metric scores from different LLM models for various datasets. The best value per dataset is shown in bold and clearly gpt-4o-mini with proportional voting outperforms all the other methods across all the evaluation measures. In this table, Llama, Mixtral and GPT refers to llama-3.1-8B, open-mistral-nemo, and gpt-4o-mini respectively. Graphical representation of this table is shown in Figure 4.
Figure 7: Venn diagram showing the overlap in the gold standard summaries obtained from three annotators (HS1, HS2 and HS3) for 5 cities – (a) City A (b) City B (c) City C (d) City D (e) City E. Each annotator selects 50 posts to be included in the summary. Note that there can be many posts with similar meaning but different annotators may choose to select different posts for gold summary.
  City A City B City C City D City E
Category HS1 HS2 HS3 HS1 HS2 HS3 HS1 HS2 HS3 HS1 HS2 HS3 HS1 HS2 HS3
 PC1 5 8 4 5 6 6 3 4 6 1 6 1 1 2 2
PC2 4 4 1 3 2 0 5 8 4 3 5 4 3 3 2
PC3 4 1 3 8 2 6 5 1 4 2 2 4 10 2 4
PC4 4 8 9 5 1 5 3 6 9 5 8 5 4 2 10
PC5 8 8 10 10 10 2 5 19 8 5 8 7 5 10 11
PC6 11 8 13 7 14 6 6 13 12 14 15 12 13 17 12
PC7 5 2 7 2 2 1 4 4 6 2 5 1 3 3 5
PC8 5 3 3 2 4 4 3 3 3 3 7 4 3 5 5
PC9 14 19 17 10 13 15 14 11 19 12 22 21 12 17 22
PC10 3 0 1 3 2 1 3 2 2 1 3 4 2 2 3
PC11 14 7 10 10 16 12 11 16 12 9 8 11 7 11 9
PC12 3 4 7 6 6 3 3 1 2 6 5 8 3 3 3
PC13 1 2 1 2 1 1 0 0 0 0 0 0 0 1 0
PC14 3 0 5 5 9 3 3 3 5 4 6 5 4 1 0
Table 10: Table demonstrates the count of various harassment cases across gold standard summaries. Note that a post can have multiple categories associated with it. Refer Table 3 for mapping of the category index (PC) with the category name.
  Posts
  One person whom my family rejected for marriage is posting my nude pictures in social media platforms, hacked my email id and forwarding the nude videos and pictures to all the contacts through different different email ids. Fake Facebook id using my pics and also fake Instagram accounts using my pics, posting bad things about me and my mother. Every day calling and texting me with different numbers. Till date he has taken 10 sim cards. Torturing my family members every day with 40 different numbers. My life has become hell. I have lost my job. Sole bread winner of the family with 2 aged parents 70+ years. Nobody is able to help me.
She was out shopping at a supermarket when she noticed a 35-40 year old man was taking her pictures/videos. Initially, she thought that it might just be a misunderstanding and the man must just be using his phone. But later got too suspicious and scary because he was following her where ever she was going. She even reported the incident to the staff members of that supermarket but before any actions were taken the man had escaped.
I was harassed at my workplace in 2015 at ABC Technology Solutions while working as a Programmer Analyst Trainee by a senior employee in the team who attempted to establish physical contact/advances several times and I was unable to react and I later complained to my reporting manager. Upon complaining to the HR and my reporting managers they claimed that they know rules pertaining to the sexual harassment act and did not take any corrective/legal action and within a few days, I was asked to give forced resignation/termination.
This incident took place around 7 30 in the night. I took a bus home after my college trip and while I was waiting to collect the ticket from the conductor he touched me in my private part in the upper part of my body. I was too confused and scared to speak out something. After a while I got a seat in the bus and was continuously yelled at for sitting crossing my legs and was threatened to be thrown out of the bus. It was late at night and I just wanted to get home safely
Table 11: Posts selected by LaMSUM are more detailed and provide a clearer description of the incident. Detailed posts enable stakeholders to gain a deeper understanding of the incident’s context, facilitating more informed and effective decision making.